# Radio-loudness statistics of quasars from Quaia–VLASS

### v1

@@ 3.5 Radio-loud fraction across the sky
In our analysis, we mostly followed Ivezić et al. (2002) and explored if radio-loud and radio-quiet QSOs are statistically consistent with each other, including differently binned subsets. To this end, they compared the normalized histograms of radio-loud and radio-quiet sources as a function of the $i$-band absolute magnitude (similar to our Fig. 10). Above all, they identified an excess of radio-quiet QSOs in the faintest bins, followed by a more abundant radio-loud population, and then again more radio-quiet sources at the peak of the distribution.
Ivezić et al. (2002) attributed these features to noise fluctuations and claimed that differences might only be caused by a few dozen sources. In our analysis, we created similar histograms (see the right panel in Fig. 10) using $G_{\rm abs}$ (limited to objects with $G<18.5$), as opposed to $i_{\rm abs}$ and an $i<18.5$ cut in previous work. Further, we analyzed more sources with 7,431 radio-loud QSOs at $G<18.5$, compared to their 280 radio-loud sources at $i<18.5$. Consequently, our larger data set is expected to help clarify the nature of the excess radio-loud populations in some $G_{\rm abs}$ bins, that were identified as random patterns in previous work (Ivezić et al., 2002).
Considering the brighter $G<18.5$ subset of our Quaia–VLASS quasars, we found that an excess of radio-loud sources does persist at $G_{\rm abs}\approx-24$, and given the statistical errors in our sample, this pattern certainly reaches several standard deviations in three neighbouring bins (see the right panel in Fig. 10).
On the left side of Fig. 10, we present an analogous histogram for the full Quaia–VLASS sample ($G<20.5$) for a more inclusive analysis. We found that the excess of radio-loud sources persists even in the full sample, albeit in a less pronounced manner. In this larger catalogue, we also noted a similar radio-loud excess for the brightest sources at $-30<G_{\rm abs}<-27$. Identifying its origin and exact significance is beyond the scope of this paper, as they might be attributed to yet-unknown systematic effects in the data, or they might be sourced by the Quaia photometric redshift errors.
@@ 4 Summary and conclusions
Driven by the availability of modern survey data sets to improve our overall understanding of the astrophysical and cosmological properties of QSOs, we cross-matched the optically selected Quaia (Gaia+WISE) catalogue of quasars with the VLASS radio sources. In our methodology, we closely followed Ivezić et al. (2002) who also analysed the properties of radio-loud QSOs in a similar context, but using significantly smaller data sets. Our main results are the following:
@@
    We created a Quaia–VLASS catalogue of 43,650 QSOs at declinations $\delta>-40\degr$ with matched radio counterparts, and we make this new catalogue publicly available for further analyses.
@@
  3. 3.
    We identified weak trends of radio-loudness changes as a function of the Quaia completeness and as a function of redshift, both pointing to possible systematic effects as sources.
@@
    We followed up on the findings by Ivezić et al. (2002) in the context of $G_{\rm abs}$ absolute magnitude statistics, and confirmed the previously seen excess of radio-loud quasars at about $G_{\rm abs}\approx-24$, for the brightest optical sources ($G<18.5$).
In summary, here we provided a detailed presentation of the Quaia–VLASS catalogue, and performed various statistical analyses in different subsets of the data to spot possible discrepancies or unexpected trends. While our findings already provided new insights on the radio properties of quasars, the upcoming release of the VLASS Epoch3 catalogue will facilitate an even more advanced analysis. With sources much fainter in the radio, we will further explore the dichotomy between radio-loud and radio-quiet sources, as well as possible relations between the positions of the QSOs in the large-scale cosmic web around them, among other potential applications.
Data availability
The Quaia quasar catalogue and its mask and selection function are publicly available [^5] https://zenodo.org/records/10403370. The VLASS radio source catalogue is also publicly available [^6] https://cirada.ca/vlasscatalogueql0, and we will make our cross-matched Quaia–VLASS data sets, codes and user instructions available publicly [^7] See http://konkoly.hu for information about our data..
@@ Acknowledgments
The Large-Scale Structure (LSS) research group at Konkoly Observatory has been supported by a Lendület excellence grant by the Hungarian Academy of Sciences (MTA). This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement number 101130774. Funding for this project was also available in part through the Hungarian National Research, Development and Innovation Office (NKFIH, grants OTKA NN147550 and K134213). S.F. was supported by the NKFIH excellence grant TKP2021-NKTA-64. L.S.-M was partially supported by the Bulgarian Ministry of Education and Science under Agreement D01-326/04.12.2023. The authors thank Kate Storey-Fisher and the Quaia team for their help with the input quasar catalogue.

### v2

@@ 3.3 Comparison to Previous Works
In this Section, we mostly followed Ivezić et al. (2002) and explored if radio-loud and radio-quiet QSOs are statistically consistent with each other, including differently binned subsets. To this end, they compared the normalized histograms of radio-loud and radio-quiet sources as a function of the $i$-band absolute magnitude (similar to our Figure 10). Above all, they identified an excess of radio-quiet QSOs in the faintest bins, followed by a more abundant radio-loud population, and then again more radio-quiet sources at the peak of the distribution.
Ivezić et al. (2002) attributed these features to noise fluctuations and claimed that differences might only be caused by a few dozen sources. In our analysis, we created similar histograms (see the right panel in Figure 10) using $G_{\rm abs}$ (limited to objects with $G<18.5$), as opposed to $i_{\rm abs}$ and an $i<18.5$ cut in their work. Further, we analyzed more sources with 7,431 radio-loud QSOs at $G<18.5$, compared to their 280 radio-loud sources at $i<18.5$. Consequently, our larger data set is expected to help clarify the nature of the excess radio-loud populations in some $G_{\rm abs}$ bins, that were identified as random patterns in their work.
Considering the brighter $G<18.5$ subset of our Quaia–VLASS quasars, we found that an excess of radio-loud sources does persist at $G_{\rm abs}\approx-24$, and given the statistical errors in our sample, this pattern certainly reaches several standard deviations in three neighbouring bins (see the right panel in Figure 10).
On the left side of Figure 10, we present an analogous histogram for the full Quaia–VLASS sample ($G<20.5$) for a more inclusive analysis. We found that the excess of radio-loud sources persists even in the full sample, albeit in a less pronounced manner. In this larger catalog, we also noted a similar radio-loud excess for the brightest sources at $30<G_{\rm abs}<-27$. Identifying its origin and exact significance is beyond the scope of this paper, as they might be attributed to yet-unknown systematic effects in the data, or they might be sourced by the Quaia photometric redshift errors.
@@ 4 Summary and conclusions
Driven by the availability of modern survey data sets to improve our overall understanding of the astrophysical and cosmological properties of QSOs, we cross-matched the optically selected Quaia (Gaia+WISE) catalog of quasars with the VLASS radio sources. In our methodology, we closely followed Ivezić et al. (2002) who also analysed the properties of radio-loud QSOs in a similar context, but using significantly smaller data sets. Our main results are the following:
@@
    We created a Quaia–VLASS catalog of 43,650 QSOs at declinations $\delta>-40\arcdeg$ with matched radio counterparts, and we make this new catalog publicly available for further analyses.
@@
    This new Quaia–VLASS catalogue provides $\approx$ 14 times more QSOs compared to Ivezić et al. (2002) across $\approx 24$ times larger sky area. Similarly, due to larger sky coverage, we provide $\approx 32,000$ optical-radio QSO cross-matches outside of the sky area of Baloković et al. (2012), and $\approx 35,000$ cross-matches outside of Arnaudova et al. (2024).
  3. 3.
@@
    We identified weak trends of radio loudness changes as a function of the Quaia selection function pointing to possible systematic effects as sources.
  5. 5.
    We followed up on the findings by Ivezić et al. (2002) in the context of $G_{\rm abs}$ absolute magnitude statistics, and confirmed with higher statistical significance the previously seen excess of radio-loud quasars at about $G_{\rm abs}\approx-24$, for the brightest optical sources ($G<18.5$).
  6. 6.
    We find that RLF increases with luminosity and decreases with redshift as opposed to previous studies in the literature, but this could be affected by a loss of completeness with increasing redshift and decreasing luminosity.
  7. 7.
    We analyse the GtG-t distribution of sources and find, given the constraints of our data, no evidence for a radio-loudness-dichotomy
In summary, here we provided a detailed presentation of the Quaia–VLASS catalog, and performed various statistical analyses in different subsets of the data to spot possible discrepancies or unexpected trends. While our findings already provided new insights on the radio properties of quasars, the upcoming release of the VLASS Epoch 3 catalog and a stacked catalog of the first three epochs (currently under construction) will facilitate an even more advanced analysis. With sources much fainter in the radio, we will further explore the dichotomy between radio-loud and radio-quiet sources, as well as possible relations between the positions of the QSOs in the large-scale cosmic web around them, among other potential applications.
5 Catalog of Quaia–VLASS cross-matches and data availability
The catalog of Quaia–VLASS cross-matches obtained in this work is depicted in Table 3, which shows column names, units, column descriptions, origin of the column and one example entry. More information about the sources can be obtained by joining them with the original catalogs via the associated identifier (Gaia, unWISE or VLASS). The cross-matches catalog is publicly available at the webpage of the online publication. It is also available with open-source code for replicating the catalog, this paper’s results, and figures at 10.5281/zenodo.16035690.
Further, the Quaia quasar catalog, its mask and selection function are made by Storey-Fisher et al. (2024) publicly available (Storey-Fisher et al., 2023) [^2] https://doi.org/10.5281/zenodo.10403370, as well as the VLASS radio source catalog [^3] https://cirada.ca/vlasscatalogueql0.
Table 3: Quaia–VLASS cross-matches catalog
Column Name Symbol Unit Description From Example Entry Value
source_id Gaia DR3 source identifier Quaia 2846730608989264640
unwise_objid unWISE DR1 source identifier Quaia 0000p212o0014594
redshift_quaia $z$ spectrophotometric redshift estimate Quaia 0.9070
[…]
ra deg barycentric R.A. of the source in ICRS at 2016.0 Quaia 0.006143428
dec deg barycentric declination of the source in ICRS at 2016.0 Quaia 21.276711143
phot_g_mean_mag $G^{\prime}$ mag Gaia $G$-band mean magnitude (not ext. corrected) Quaia 20.197878
phot_bp_mean_mag mag Gaia integrated BP mean magnitude Quaia 20.221630
phot_rp_mean_mag mag Gaia integrated RP mean magnitude Quaia 19.663361
mag_w1_vg mag unWISE W1 magnitude Quaia 15.583867
mag_w2_vg mag unWISE W2 magnitude Quaia 14.950078
Component_name unique VLASS name VLASS VLASS2QLCIR
J000001.47+211636.0
RA deg right ascension in VLASS VLASS 0.00614736
DEC deg declination in VLASS VLASS 21.276691
Total_flux $F_{\mathrm{tot}}$ mJy integrated radio flux density VLASS 1.846
reddening $r$ sky area’s reddening (Eq. 2), from Schlegel et al. (1998) this work 0.052644
t $t$ mag radio magnitude (Eq. 3) this work 15.734417
g $G$ mag extinction corrected $G$-band magnitude (Eq. 2) this work 20.053634
g_abs $G_{\mathrm{abs}}$ mag absolute extinction corrected $G$-band magnitude (Eq. 5) this work 23.845170-23.845170
sep $\Delta\theta$ arcsec angular separation between VLASS and Quaia coordinates this work 0.07372
r_loudness radio-loud (1) or radio-quiet (0) (Eq. 4) this work 1
@@ Acknowledgments
The authors thank the anonymous referee for their helpful and constructive suggestions, which improved the clarity and quality of this manuscript. Further thanks are due to Kate Storey-Fisher and the Quaia team for their feedback and assistance with the input quasar catalog. The Large-Scale Structure (LSS) research group at Konkoly Observatory has been supported by a Lendület excellence grant by the Hungarian Academy of Sciences (MTA). This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement number 101130774. Funding for this project was also available in part through the Hungarian National Research, Development and Innovation Office (NKFIH, grants OTKA NN147550 and K134213). S.F. was supported by the NKFIH excellence grant TKP2021-NKTA-64. L.S.-M. was partially supported by the Bulgarian Ministry of Education and Science under Agreement D01-326/04.12.2023.
ORCID IDs
Nestor Arsenov: 0009-0009-8592-5400
Sándor Frey: 0000-0003-3079-1889
András Kovács: 0000-0002-5825-579X
Lyuba Slavcheva-Mihova: 0000-0002-1582-4913